Please assist to explain how the following can be proved by induction:

1+3+9+ ... + 3n-1 = (3 – 1)/2
Answer #1

Q :- 1+3+9+ ... + 39.13 (31-1272 --) Box proove above by induction we have pooove thyce PC12,Pck) and PC12+12 For Paul n=1 1For -PCKEL2_6_nektl (+3+9+...+ 3.17-1 +3 k = Get alle Take R.H.S= 1 +3+9 taas +371a/ tak K (3 -1 t3k 2 3R_/ +2.3k 2 a 3. 3* +